suppose that we have a qualitative variable month with categories January February etc how many dummy variables are needed to escribe month
Answer #1

Answer

We know that there are 12 months in a year

Number of dummy variables is given as

dummy variables = k - 1

where k is the total number of variables

In this case, we have 12 variables for 12 months

so, number of dummy variables = k-1

= 12-1

= 11